Explain why the 0.20 mol L -1 sodium hydroxide solution could not have been prepared by weighing out 2.00g sodium hydroxide, dissolving this in water and making up to 500 ml.
Answer #1

Mass of sodium hydroxide = 2 gram

Molar mass of NaOH = 40 g/mol

So moles = mass÷molar mass = 2/40 = 0.05 moles

Volume of solution = 500 ml = 0.5 L

So molarity = moles/volume = 0.05/0.5 = 0.1 mol/L

As we are getting molarity as 0.1 M

Hence, we can't prepare 0.2 mol/L

Amount of NaOH required is less